MUC-NOTE-COOL THEORY-xxx

Status of Neutrino Factory and Muon Collider Research and Development and Future Plans

Mohammad M. Alsharo’a,^{1} Charles M. Ankenbrandt,^{2} Bruno, R. Autin,^{3} Valeri I. Balbekov,^{2} Vernon D. Barger,^{4} Odette Benary,^{5} J. Roger J. Bennett,^{6} J. Scott Berg,^{7} Mike

S. Berger,^{8} Martin Berz,^{9} Edgar L. Black,^{1} S. Alex Bogacz,^{10} Stephen B. Bracker,^{11} Alan David Bross,^{2} Luca Bruno,^{3} Allen C. Caldwell,^{12} Kevin W. Cassel,^{1} David B. Cline,^{13} Linda R. Coney,^{12} John N Corlett,^{14} Lucien M. Cremaldi,^{11} mary anne c.

cummings,^{15} Christine Darve,^{2} Fritz DeJongh,^{2} Paul Drumm,^{6} V. Daniel Elvira,^{2} Deborah Errede,^{16} Adrian Fabich,^{3} William M. Fawley,^{14} Richard C. Fernow,^{7} Massimo Ferrario,^{17}

Nathaniel J. Fisch,^{18} Yasuo Fukui,^{13} Tony A. Gabriel,^{19} Raphael Galea,^{12} Juan C. Gallardo,^{7} Roland Garoby,^{3} Steve Geer,^{2} Simone Gilardoni,^{3} Romulus Godang,^{11}

Maury Goodman,^{20} Michael R. Gosz,^{1} Michael A. Green,^{14} John F. Gunion,^{21} Ramesh Gupta,^{7} John R Haines,^{19} Gail G. Hanson,^{22} Don Hartill,^{23} Helmut D. Haseroth,^{3} Kara Hoffman,^{24} Rolland P. Johnson,^{1} Stephen A. Kahn,^{7} Daniel M. Kaplan,^{1} Eberhard K. Keil,^{2} Bruce J. King,^{25} Harold G. Kirk,^{7} Yoshitaka Kuno,^{26} Tony S. Ladran,^{14} Wing W Lau,^{27} Jacques A. Lettry,^{3} Derun Li,^{14} Changguo Lu,^{18} Kyoko Makino,^{16} Kirk T McDonald,^{18} Frederick E. Mills,^{2} Nikolai V. Mokhov,^{2} Al Moretti,^{2}

Jim Norem,^{20} Yasar Onel,^{28} Mark J. Oreglia,^{24} Robert B Palmer,^{7} Brett Parker,^{7} Zohreh Parsa,^{7} Gregory Penn,^{29} Milorad B Popovic,^{2} zubao qian,^{2} Rajendran

Raja,^{2} Helge Leif Ravn,^{3} Pavel Rehak,^{7} Robert A. Rimmer,^{14} Thomas J. Roberts,^{1} Thomas Roser,^{7} Robert Rossmanith,^{30} Roman V. Samulyak,^{7} Stefan Schlenstedt,^{12} Peter Schwandt,^{8} Andrew M. Sessler,^{14} Peter Sievers,^{3} Gregory I. Silvestrov,^{31} Nick

Simos,^{7} Alexander N. Skrinsky,^{31} Philip Thomas Spampinato,^{19} Panagiotis Spentzouris,^{2} Raymond Joseph Stefanski,^{2} Peter Stoltz,^{32} Sergei I. Striganov,^{28} Iuliu Stumer,^{7} Donald J.

Summers,^{11} Lee C. Teng,^{20} Peter A. Thieberger,^{7} maury tigner,^{23} Michael Todosow,^{7} Yagmur Torun,^{1} Dejan Trbojevic,^{7} Zafar U. Usubov,^{2} Andreas J. Van Ginneken,^{2} Tatiana A. Vsevolozhskaya,^{31} Chun-xi Wang,^{20} Robert J. Weggel,^{7} Erich H. Willen,^{7} Jonathan S. Wurtele,^{21} Moohyun Yoon,^{33} Yongxiang Zhao,^{7} and Michael S. Zisman^{14}

1Illinois Institute of Technology 2Fermilab 3CERN

4University of Wisconsin 5Tel Aviv University / BNL

6Rutherford Appleton Laboratory 7Brookhaven National Laboratory

8Indiana University 9Michigan State University

10Thomas Jefferson National Accelerator Facility

11University of Mississippi

12Columbia University 13University of California, Los Angeles 14Lawrence Berkeley National Laboratory

15Northern Illinois University 16University of Illinois, Urbana-Champaign

17INFN-LNF 18Princeton University

19Oak Ridge National Laboratory 20Argonne National Laboratory 21University of California, Davis 22University of California, Riverside

23Cornell University 24University of Chicago 25Northwestern University

26Osaka University 27Oxford University 28University of Iowa

29University of California, Berkeley

30Forschungszentrum Karlsruhe 31Budker Institute of Nuclear Physics

32Tech-X Corporation

33Pohang University of Science and Technology

(Dated: May 24, 2002)

Abstract

We describe the status of the effort to realize a first Neutrino Factory and the progress made in understanding the problems associated with the collection and cooling of muons towards that end. . . . ..

INTRODUCTION

4